# Spatially-guided Temporal Aggregation for Robust Event-RGB Optical Flow Estimation

### v1

@@ V-A Datasets and Setup
To ensure fair comparisons with prior methods, we conducted extensive experiments on MVSEC [48] and DSEC-Flow [49]. Both datasets provide frame and event data from real-world scenes and are widely adopted for evaluating event-based optical flow. MVSEC includes one daytime outdoor driving sequence, one nighttime outdoor driving sequence, and four indoor flying sequences. We adhered to the same experimental setup as in previous work [35]. Specifically, Experiments on MVSEC were performed at different time intervals, $dt=1$ and $dt=4$ images. For data partitioning, outdoor_day_2 was used for training, while outdoor_day_1 and three indoor sequences formed the test set. Consistent with prior methods [35], only 800 frames from outdoor_day_1 were used for evaluation.
DSEC-Flow covers a broader range of driving scenarios, including challenging conditions such as nighttime, sunrise, sunset, and tunnels. The dataset provides an official training set and a public online benchmark. The test set of DSEC-Flow also contains various challenging scenes, making it a widely used benchmark for evaluating event-based algorithms. In line with previous studies, we submitted our results to its public benchmark for evaluation.
Metrics. The primary accuracy metric for both DSEC-Flow and MVSEC is End-Point-Error (EPE). DSEC-Flow additionally reports Angular Error (AE) and the percentage of EPE exceeding $N$ pixels (NPE). For MVSEC, we followed prior work to compute outliers, defined as predictions with an EPE greater than 3 pixels or more than 5% of the ground truth magnitude.
Train details. The proposed method was implemented using PyTorch. For event voxelization, following previous methods [26, 22], events between two frames were divided into 15 bins. The number of targets $N$ was set to 5, resulting in $\Delta t=0.2(T_{k+1}-T_{k})$. Consequently, the voxel grid $V_{T_{k}-\Delta t}^{T_{k+1}}$ consists of $18$ bins, including one reference and five targets. The voxel normalization parameter $\epsilon$ in ICE was set to 0.1. The search radius for all correlation volumes was set to 4. The decay factor $\gamma$ was set to 0.85. During training, the batch size was set to 6 and the learning rate was 0.0002. The number of iterations was fixed at 6 for both training and testing. The network was trained on DSEC for 200k steps and MVSEC for 100k steps.
@@ V-B Results of the DSEC-Flow Dataset
Accuracy. The overall evaluation results on the DSEC-Flow dataset are summarized in Table I, with detailed results for each sequence provided in Table II. Several conclusions can be drawn from these results. The methods are categorized into three groups based on input type: frame-only, event-only, and cross-modal. Both cross-modal methods outperform the state-of-the-art event-based method ECDDP [50], highlighting that incorporating frame data significantly enhances prediction accuracy. The event branches in our method and BFlow are both based on temporally dense correlation volumes, similar to TMA. To evaluate the effectiveness of the proposed fusion strategy, we compare the performance of our method and BFlow against TMA. Our method improves inter-modal interactions by replacing the simple concatenation mechanism in BFlow with cross-modal guided aggregation. Compared to TMA, BFlow reduces the EPE from 0.74 to 0.69, and our method further reduces the EPE to 0.66, with a 10% improvement in accuracy. These results validate the effectiveness of the guided fusion strategy. Qualitative results in Figs. 1 and 6 show that single-modal approaches often struggle with ambiguity in certain regions due to insufficient intensity and texture information. Our proposed cross-modal guidance strategy effectively mitigates this uncertainty, leading to more stable and consistent predictions across consecutive frames.
For reference, we also report the performance of frame-only methods. The DSEC-Flow test set contains numerous low-light scenes, which limits the performance gains of frame-based methods compared to event-based approaches. The results indicate that integrating event data can enhance the accuracy of frame-only methods, particularly in challenging environments. Finally, we compare our method to other cross-modal approaches. Compared to the state-of-the-art cross-modal method BFlow, our approach achieves a 4% reduction in EPE, demonstrating that the proposed fusion strategy retains the strengths of each modality while enhancing cross-modal collaboration. Our method also shows notable improvements over BFlow in the NPE and AE metrics. As shown in Table II, our method maintains consistent performance across different sequences, underscoring its robustness.
This study also investigates the effect of pretraining with synthetic data [23] on performance. The synthetic dataset, similar to FlyingChairs [27], differs significantly from real-world driving scenarios [26]. The network was initialized with parameters trained on this synthetic dataset, while all other settings unchanged. As shown in Table II, pretraining substantially improves network performance, highlighting the strong potential of synthetic data for enhancing optical flow accuracy.
TABLE I: DSEC-Flow evaluation results. Bold indicates the best accuracy, while underlining indicates the second best. ‘E’ represents events, and ‘I’ represents images. ‘*’ indicates that the network is pre-trained with synthetic data. $\downarrow$” indicates the smaller, the better.
@@
TABLE II: Detail results on DSEC-Flow. Bold indicates the best accuracy. All results are available on the public online benchmark of DSEC-Flow.
Method Overall interlaken_00_b interlaken_01_a thun_01_a
ERAFT [26] 0.79 2.68 2.85 1.39 6.19 2.36 0.90 3.91 2.54 0.65 1.87 2.94
TMA [22] 0.74 2.30 2.68 1.39 5.79 2.157 0.81 3.11 2.23 0.62 1.61 2.88
IDNet [51] 0.70 1.96 2.58 1.25 4.35 2.11 0.77 2.60 2.25 0.57 1.47 2.66
ECDDP [50] 0.72 2.04 2.72 1.31 5.03 2.00 0.76 2.47 2.18 0.52 1.36 2.34
Ours 0.66 1.68 2.37 1.22 4.32 1.96 0.72 2.10 2.06 0.55 1.31 2.43
Method thun_01_b zurich_city_12_a zurich_city_14_c zurich_city_15_a
ERAFT [26] 0.58 1.52 2.20 0.61 1.06 4.50 0.71 1.91 3.43 0.59 1.30 2.55
TMA [22] 0.55 1.31 2.10 0.57 0.87 4.38 0.66 1.99 3.09 0.55 1.08 2.51
IDNet [51] 0.55 1.35 2.07 0.60 1.16 4.56 0.76 2.74 3.74 0.55 1.02 2.55
ECDDP [50] 0.51 1.21 1.93 0.55 0.76 4.35 0.69 2.39 3.22 0.52 0.89 2.41
Ours 0.50 1.12 1.76 0.50 0.54 3.80 0.55 0.94 2.47 0.53 0.92 2.27
Robustness Analysis. The ICE representation integrates both events and frames, providing robust guidance across diverse scenarios. We validate the robustness of our approach by evaluating it on the DSEC-Flow test set, which contains challenging scenes. The detailed evaluation results are presented in Table II. Since comprehensive results for BFlow are unavailable, we compare our method with TMA [22] and ECDDP [50]. Additionally, Fig. 6 provides a qualitative comparison, showcasing the nighttime sequence zurich_city_12_a and the daytime sequence zurich_city_14_c as representative examples.
Figure 6: Qualitative results on DSEC-Flow. The first two rows display the night scene zurich_city_12_a with contrast-adjusted images for improved visibility. The bottom row shows the daytime scene zurich_city_14_c. The third and fourth columns compare the event-only method TMA [22] with our method.
As shown in Table II, our method achieves the highest prediction accuracy across both challenging and regular scenes. By leveraging the texture information from frames, the proposed approach demonstrates significant improvements in regular scenes. For example, our method outperforms TMA by 17% on the zurich_city_14_c sequence. The bottom row of Fig. 6 highlights that even in well-lit scenes, events may still lack sufficient texture information, which leads to errors. Our approach addresses this by using stable ICE features to compensate for the sparse event features, substantially improving accuracy. Notably, our method surpasses the state-of-the-art event-based method ECDDP [50] in nighttime scenarios, despite reduced frame quality. For instance, our method achieves a 12% accuracy improvement over ECDDP on the zurich_city_12_a sequence. This performance gain likely stems from the tendency of event cameras to generate increased noise in low-light conditions, while prominent edges and illuminated areas in frames still provide valuable information. ICE features enhance the robustness of guiding features by effectively integrating event data. Another demonstration of the robustness of our method is the prediction consistency shown in Fig. 1. Our guided fusion strategy ensures stable spatial features, providing a critical advantage over other cross-modal approaches.
@@ V-C Results of the MVSEC Dataset
We evaluated the generalization performance of our method on the MVSEC dataset, as reported in Table III. In accordance with the experimental setup, all models were trained on the outdoor_day_2 sequence and tested on the indoor and outdoor_day_1 sequences. The domain gap between outdoor_day_1 and outdoor_day_2 is relatively small, while the gap between the indoor sequence and outdoor_day_2 is more substantial. As shown in Table III, our method achieves significantly higher accuracy on the indoor sequence compared to event-only approaches, while performing less effectively on the outdoor_day_1 sequence. These results indicate that our fusion strategy reduces overfitting, thereby improving generalization.
TABLE III: Evaluation results on MVSEC. Bold indicates the best accuracy. ‘%Outlier’ denotes the proportion of outliers to all valid pixels.
Method Input indoor_flying_1 indoor_flying_2
dt=1 dt=4 dt=1 dt=4
EPE($\downarrow$) %Outlier($\downarrow$) EPE($\downarrow$) %Outlier($\downarrow$) EPE($\downarrow$) %Outlier($\downarrow$) EPE($\downarrow$) %Outlier($\downarrow$)
ERAFT [26] E 1.10 5.72 2.81 40.25 1.94 30.79 5.09 64.19
TMA [22] E 1.06 3.63 2.43 29.91 1.81 27.29 4.32 52.74
Ours E+I 0.89 2.38 2.27 28.03 1.55 18.49 4.16 48.38
Method Input indoor_flying_3 outdoor_day_1
dt=1 dt=4 dt=1 dt=4
EPE($\downarrow$) %Outlier($\downarrow$) EPE($\downarrow$) %Outlier($\downarrow$) EPE($\downarrow$) %Outlier($\downarrow$) EPE($\downarrow$) %Outlier($\downarrow$)
ERAFT [26] E 1.66 25.20 4.46 57.11 0.24 0.00 0.72 1.12
TMA [22] E 1.58 23.26 3.60 42.02 0.25 0.07 0.70 1.08
Ours E+I 1.19 13.22 3.01 33.24 0.32 0.00 0.79 2.04
Figure 7: Qualitative ablation studies. When significant overlap occurs between foreground and background events, the network tends to experience regional confusion. Introducing frame data alone does not fully address this issue. Our proposed cross-modal guided aggregation strategy reduces errors by using stable ICE features to guide event features. Additionally, replacing single-modal context with compact spatiotemporal context features leads to notable performance improvements.
@@ V-D Ablation Study
Spatiotemporal Context. We introduce the mix-fusion context encoder to extract spatiotemporal context features, in place of single-modality features. Previous studies have typically extracted either temporal context features from event voxels or spatial context features from images. To investigate the differences between these context features, we replace the fused spatiotemporal features with frame context features and event context features, respectively. For fair comparison, the spatiotemporal features output by MFCE are dimensionally matched to the single-modality features. As shown in rows 3 and 5 of Table IV, extracting context features solely from events results in reduced accuracy (EPE increases from 0.66 to 0.68), highlighting the advantage of spatiotemporal features over event-only features. In row 2, following the typical settings of frame-based optical flow methods, we extract context features exclusively from the first frame. Compared to spatiotemporal features, spatial features lead to a notable drop in accuracy (EPE increases from 0.66 to 0.69), likely due to the lack of rich motion cues available in event data. Furthermore, comparing event context features with frame context features reveals that event features outperform frame-based features. This may be attributed to events capturing both motion cues from intermediate processes and the spatial structure of the scene.
To further validate the effectiveness, we conducted a qualitative comparison. As shown in Fig. 7, the single modality context feature is limited in information, leading to blurred boundaries between the guardrail and background. In contrast, the spatiotemporal context feature delivers clearer boundary predictions by leveraging complementary modalities. In conclusion, although both types of features could provide context information during the iterative refinement process, event-based context features typically yield better performance. The spatiotemporal context features further enhance this by offering more comprehensive information and compact representations.
Cross-Modal Guided Aggregation. To enhance inter-modal interactions and achieve a deeper fusion of events and frames, we proposed a cross-modal guided aggregation strategy. Unlike BFlow, which concatenates motion features from both modalities directly, our approach leverages the spatial stability of ICE motion features to guide the aggregation of temporally-dense event motion features. The aggregation module is built on a transformer architecture, where ICE features serve as keys and values to guide and enrich the spatially sparse event motion features. In the ablation study, we replace guided aggregation with direct concatenation for comparison. As shown in rows 4 and 5 of Table IV, the guided aggregation strategy significantly improves the accuracy, confirming that enhancing inter-modal interactions improves modality fusion compared to simple concatenation.
To further validate its effectiveness, we present a qualitative comparison. As illustrated in Fig. 7, while spatiotemporal context features help delineate the boundaries between the foreground and background, background motion shows prediction errors caused by occlusion from the guardrail. Our proposed method mitigates these errors through stable ICE guidance, demonstrating the effectiveness of the guided fusion strategy.
@@ V-E Efficiency
Figure 8: EPE vs. Number of iterations at inference. The figure illustrates prediction accuracy across iterations during inference. Our method exceeds the accuracy of TMA-6 in just two iterations and achieves the accuracy of BFlow-12 in only three iterations, where ‘-N’ indicates the number of iterations.
@@
Methods Params Inference time(ms) iterations EPE
ERAFT [26] 5.3M 74 12 0.79
TMA [22] 6.9M 43 6 0.74
BFlow [23] 6.5M 111 12 0.69
Ours 9.2M 61 6 0.66
@@
We present the inference time and performance of different approaches in Table V. For correlation-based methods, inference time is significantly affected by the number of iterations [52]. As shown in the table, inference time consistently increases with the number of iterations across different methods, revealing a clear trade-off between speed and accuracy. Reducing iterations accelerates inference but leads to a notable drop in accuracy. Our approach addresses this by enhancing the quality of extracted features, enabling higher accuracy with fewer iterations.
We first compare the inference time of our method with TMA to evaluate the overhead from incorporating additional modalities. With the same number of iterations, incorporating new modalities increases inference time by just 18 ms compared to TMA. Despite the slight increase in inference time, the 10% accuracy gain over TMA makes the introduction of new modalities highly beneficial. Next, we compare inference time with other cross-modal methods to assess the efficiency of different fusion strategies. Our method outperforms the state-of-the-art fusion-based method [23] with only half the number of iterations (6 vs. 12), indicating that our aggregated features encode richer and more precise motion cues for refinement. Our method reduces inference time by 45% and lowers end-point error by 4% compared to BFlow. Overall, our approach delivers superior performance in the early iterations. As shown in Fig. 8, our method surpasses the accuracy of TMA-6 at the second iteration and achieves the accuracy of BFlow-12 by the third iteration.
VI Conclusion and limitation

### v2

@@ V Experiments
TABLE I: Detail results on DSEC-Flow. Best results are highlighted as first, and second. ‘E’ represents events, and ‘I’ represents images. $\downarrow$” indicates the smaller, the better. All results are available on the public online benchmark of DSEC-Flow.
Method Input Overall interlaken_00_b interlaken_01_a thun_01_a
ERAFT [26] E 0.79 2.68 2.85 1.39 6.19 2.36 0.90 3.91 2.54 0.65 1.87 2.94
TMA [22] E 0.74 2.30 2.68 1.39 5.79 2.157 0.81 3.11 2.23 0.62 1.61 2.88
IDNet [49] E 0.70 1.96 2.58 1.25 4.35 2.11 0.77 2.60 2.25 0.57 1.47 2.66
ECDDP [50] E 0.72 2.04 2.72 1.31 5.03 2.00 0.76 2.47 2.18 0.52 1.36 2.34
BFlow [23] E+I 0.69 2.02 2.42 1.11 4.58 1.93 0.77 2.63 2.07 0.59 1.59 2.71
Ours E+I 0.66 1.68 2.37 1.22 4.32 1.96 0.72 2.10 2.06 0.55 1.31 2.43
Method Input thun_01_b zurich_city_12_a zurich_city_14_c zurich_city_15_a
ERAFT [26] E 0.58 1.52 2.20 0.61 1.06 4.50 0.71 1.91 3.43 0.59 1.30 2.55
TMA [22] E 0.55 1.31 2.10 0.57 0.87 4.38 0.66 1.99 3.09 0.55 1.08 2.51
IDNet [49] E 0.55 1.35 2.07 0.60 1.16 4.56 0.76 2.74 3.74 0.55 1.02 2.55
ECDDP [50] E 0.51 1.21 1.93 0.55 0.76 4.35 0.69 2.39 3.22 0.52 0.89 2.41
BFlow [23] E+I 0.55 1.42 1.84 0.54 0.72 3.86 0.59 1.37 2.44 0.57 1.24 2.38
Ours E+I 0.50 1.12 1.76 0.50 0.54 3.80 0.55 0.94 2.47 0.53 0.92 2.27
@@ V-A Datasets and Setup
To ensure fair comparisons with prior methods, we conducted extensive experiments on MVSEC [51] and DSEC-Flow [52]. Both datasets provide frame and event data from real-world scenes and are widely adopted for evaluating event-based optical flow. MVSEC includes outdoor driving sequences and several indoor flying sequences. We followed the same setup as in prior work [36], which set time intervals to $dt=1$ and $dt=4$ images. Generally, outdoor_day_2 was used for training, while outdoor_day_1 and three indoor sequences formed the test set. DSEC-Flow covers a broader range of driving scenarios, including challenging conditions such as nighttime, sunrise, sunset, and tunnels. The dataset provides an official training set and a public online benchmark. The test set of DSEC-Flow also contains various challenging scenes, making it a widely used benchmark for evaluating event-based algorithms. In line with previous studies, we submitted our results to its public benchmark for evaluation.
Metrics. The primary accuracy metric for both DSEC-Flow and MVSEC is End-Point-Error (EPE). DSEC-Flow additionally reports Angular Error (AE) and the percentage of EPE exceeding $N$ pixels (NPE). For MVSEC, we followed prior work to compute outliers, defined as predictions with an EPE greater than 3 pixels or 5% of the ground truth.
Train Details. The proposed method was implemented using PyTorch. For event voxelization, following previous methods [26, 22], events between two frames were divided into 15 bins. The number of targets $N$ was set to 5, resulting in $\Delta t=0.2(T_{k+1}-T_{k})$. The voxel normalization parameter $\epsilon$ in ICE was set to 0.1. The radius for lookup was set to 4. The decay factor $\gamma$ was set to 0.85. During training, the batch size was set to 6 and the learning rate was 0.0002. The number of iterations was fixed at 6 for both training and testing. The network was trained on DSEC for 200k steps and MVSEC for 100k steps.
@@ V-B Results of the DSEC-Flow Dataset
Accuracy. Detailed per-sequence results on DSEC-Flow are shown in Table I, and additional results are provided in Table II. Several conclusions can be drawn from these results. Both cross-modal methods outperform the state-of-the-art event-based method ECDDP [50], highlighting that incorporating frame data significantly enhances prediction accuracy. The event-only method TMA [22], fusion-based method BFlow [23], and our method are all based on temporally dense correlation volumes. Among these approaches, our method improves inter-modal interactions by replacing the simple concatenation mechanism in BFlow with cross-modal guided aggregation. Our method also shows notable improvements over BFlow in the NPE and AE metrics, maintains consistent performance across different sequences, underscoring its robustness. Qualitative results in Figs. 1 and 4 show that the proposed cross-modal guidance strategy effectively mitigates the ambiguity of flow, leading to more stable and consistent predictions across consecutive frames. Additionally, in Fig. II, our method outperforms both single-modal and cross-modal methods, while pretraining on synthetic data [23] further improves the performance.
TABLE II: DSEC-Flow evaluation results. Best results are highlighted as first, second, and third. ‘*’ indicates pre-trained with synthetic data.
@@
Robustness Analysis. The ICE representation integrates both events and frames, providing robust guidance across diverse scenarios. We validate the robustness of our approach by evaluating it on the DSEC-Flow test set, which contains challenging scenes. The detailed evaluation results are presented in Table I. Additionally, Fig. 4 provides a qualitative comparison, showcasing the nighttime sequence zurich_city_12_a and the snowy sequence interlaken_01_a as examples.
Figure 4: Qualitative results on DSEC-Flow. The night and daytime sequences are from zurich_city_12_a and interlaken_01_a, respectively. Our method outperforms both the event-only TMA [22] and the best fusion-based method BFlow [23].
As shown in Table I, our method achieves the best accuracy across both challenging and regular scenes. Notably, our method surpasses the state-of-the-art event-based method ECDDP [50] in nighttime scenarios, despite reduced frame quality. For instance, our method achieves a 12% accuracy improvement over ECDDP on the zurich_city_12_a sequence. This performance gain likely stems from the tendency of event cameras to generate increased noise in low-light conditions, while prominent edges and illuminated areas in frames still provide valuable information. ICE features enhance the robustness of guiding features by effectively integrating event data. Another demonstration of the robustness of our method is the prediction consistency shown in Fig. 1. Our guided fusion strategy ensures stable spatial features, providing a critical advantage over other cross-modal approaches.
@@ V-C Results of the MVSEC Dataset
We evaluated the generalization performance on the MVSEC dataset, as reported in Table III. In accordance with the experimental setup, all models were trained on the outdoor_day_2 sequence and tested on the indoor and outdoor_day_1 sequences. The domain gap between outdoor_day_1 and outdoor_day_2 is relatively small.
From the results, we observe that model-based and unsupervised methods generally exhibit stronger generalization. In particular, MultiCM [34] achieves the best performance across all indoor sequences without any training or labeled data. This highlights the potential overfitting associated with supervised learning in domain-shift scenarios.
Interestingly, our method outperforms other supervised event-only methods on the indoor sequence but performs relatively worse on outdoor_day_1. We hypothesize that this contrast reflects the regularizing effect of incorporating image-based guidance, which helps mitigate overfitting by providing complementary spatial information. These results support the effectiveness of our fusion strategy in improving generalization to unseen domains.
TABLE III: Evaluation results on MVSEC. Best results are highlighted as first, second, and third. ‘%Outlier’ denotes the proportion of outliers to all valid pixels.
Method Input indoor_flying_1 indoor_flying_2 indoor_flying_3 outdoor_day_1
EPE %Outlier EPE %Outlier EPE %Outlier EPE %Outlier
dt=1
MB Nagata et al. [53] E 0.62 0.93 0.84 0.77
Brebion et al. [54] E 0.52 0.10 0.98 5.50 0.71 2.10 0.53 0.20
MultiCM(Burgers’) [34] E 0.42 0.10 0.60 0.59 0.50 0.28 0.30 0.10
USL EV-FlowNet [55] E 0.58 0.00 1.02 4.00 0.87 3.00 0.32 0.00
FireFlowNet [17] E 0.97 2.60 1.67 15.30 1.43 11.00 1.06 6.60
ConvGRU-EV-FlowNet [37] E 0.60 0.51 1.17 8.06 0.93 5.64 0.47 0.25
SSL Ev-FlowNet [36] E 1.03 2.20 1.72 15.10 1.53 11.90 0.49 0.20
Spike-FlowNet [56] E 0.84 1.28 1.11 0.49
Ziluo et al. [57] E 0.57 0.10 0.79 1.60 0.72 1.30 0.42 0.00
SL ERAFT [26] E 1.10 5.72 1.94 30.79 1.66 25.20 0.24 0.00
TMA [22] E 1.06 3.63 1.81 27.29 1.58 23.26 0.25 0.07
Ours E+I 0.89 2.38 1.55 18.49 1.19 13.22 0.32 0.00
dt=4
MB MultiCM(Burgers’) [34] E 1.69 12.95 2.49 26.35 2.06 19.03 1.25 9.21
USL EV-FlowNet [55] E 2.18 24.20 3.85 46.80 3.18 47.80 1.30 9.70
ConvGRU-EV-FlowNet [37] E 2.16 21.51 3.90 40.72 3.00 29.60 1.69 12.50
SSL Ev-FlowNet [36] E 2.25 24.70 4.05 45.30 3.45 39.70 1.23 7.30
Spike-FlowNet [56] E 2.24 3.83 3.18 1.09
Ziluo et al. [57] E 1.77 14.70 2.52 26.10 2.23 22.10 0.99 3.90
SL ERAFT [26] E 2.81 40.25 5.09 64.19 4.46 57.11 0.72 1.12
TMA [22] E 2.43 29.91 4.32 52.74 3.60 42.02 0.70 1.08
Ours E+I 2.27 28.03 4.16 48.38 3.01 33.24 0.79 2.04
Figure 5: Qualitative ablation studies. Foreground–background overlap causes regional confusion. Our method mitigates these errors, highlighting the contribution of each individual component.
@@ V-D Ablation Study
Spatiotemporal Context. We introduce the mix-fusion context encoder to extract spatiotemporal context features, in place of single-modality features. To investigate the differences between different context features, we replace the spatiotemporal features with frame or event context features, respectively. As shown in rows 2 and 3 of Table IV, extracting context features solely from events or frame results in reduced accuracy, highlighting the advantage of spatiotemporal features over single-modality features. The qualitative comparison in Fig. 5 further validate the effectiveness. Furthermore, comparing event context features with frame context features reveals that event features outperform frame-based features. This may be attributed to events capturing both motion cues from intermediate processes and the spatial structure of the scene. In conclusion, the spatiotemporal context features offer more comprehensive information and compact representations.
Cross-Modal Guided Aggregation. Different with BFlow, which concatenates motion features from both modalities directly, our approach leverages the spatial stability of ICE motion features to guide the aggregation of temporally dense event motion features. In the ablation study, we replace guided aggregation with direct concatenation for comparison. As shown in rows 4 and 5 of Table IV, the guided aggregation strategy significantly improves the accuracy, confirming that enhancing inter-modal interactions improves modality fusion compared to simple concatenation. Fig. 5 further validates this conclusion through qualitative comparisons.
@@ V-E Efficiency
Figure 6: EPE vs. inference iterations. Our method achieves faster convergence, surpassing TMA-6 in 2 iterations and reaching BFlow-12 accuracy in just 3 iterations.
@@
Methods Params iterations Times(ms) Mem.[MB] EPE
ERAFT [26] 5.3M 12 74 975 0.79
TMA [22] 6.9M 6 43 2185 0.74
BFlow [23] 6.5M 12 111 1887 0.69
Ours 9.2M 6 61 2459 0.66
@@
We present the iteration efficiency and computational performance of different approaches in Fig. 6 and Table V. For correlation-based methods, inference time is significantly affected by the number of iterations [58]. As shown in the table, inference time consistently increases with the number of iterations across different methods, revealing a clear trade-off between speed and accuracy. Reducing iterations accelerates inference but leads to a notable drop in accuracy. Our approach addresses this by enhancing the quality of extracted features, enabling higher accuracy with fewer iterations. As shown in Fig. 6, our approach delivers superior performance in the early iterations, which surpasses the accuracy of TMA-6 at the second iteration and achieves the accuracy of BFlow-12 by the third iteration.
We further report the GPU memory consumption. As shown in Table V, multi-correlation volume methods (TMA, BFlow, and ours) naturally incur higher memory usage compared to single-volume baselines. Our method exhibits slightly higher memory usage than TMA, consistent with the additional frame modality, while still maintaining high inference efficiency and superior accuracy.
V-F Downstream Task
Optical flow provides dense correspondence between frames and serves as a strong initialization for frame interpolation. To isolate the role of flow, we warp the end frame backward using the predicted flow and compare the result with the start frame using SSIM and PSNR. This setup allows us to indirectly assess the interpolation capability of the flow. Quantitative results are reported in Table VI, and qualitative comparisons are shown in Fig. 7. The results demonstrate that: (i). Predicted flow enables the reconstruction of most pixels with reasonable accuracy, validating its utility in interpolation. (ii). Our method achieves higher accuracy around object boundaries and fine structures, along with improved overall similarity.
Figure 7: Qualitative comparison of warped frames obtained using different optical flow methods.
TABLE VI: Evaluation of flow accuracy by warping frame on the DSEC test. SSIM and PSNR are computed between the original and flow warped frame.
SSIM/PSNR inter_01_a inter_00_b thun_01_a city_12_a
BFlow 0.788/25.14 0.675/18.78 0.793/25.20 0.826/28.11
Ours 0.792/25.14 0.680/18.90 0.804/25.53 0.836/28.52
city_15_a city_14_c thun_01_b Overall
BFlow 0.770/22.36 0.802/23.50 0.782/24.10 0.776/24.12
Ours 0.772/22.45 0.802/23.52 0.785/24.14 0.781/24.24
Figure 8: Failure case under extreme low-light conditions where both RGB and event inputs are severely degraded.
V-G Limitations and Future Work
While our method demonstrates strong performance across diverse scenarios, it has limitations in extreme cases where both RGB and event inputs are severely degraded. A typical failure case is observed under extreme low-light conditions, where the RGB input is nearly black and event data are sparse and noisy. As shown in Fig. 8, the ICE representation fails to provide spatially stable guidance, resulting in degraded flow estimation. A potential improvement involves aggregating events across longer durations with motion compensation to reconstruct spatially structures, which may provide alternative guidance in the absence of reliable RGB input.
VI Conclusion